# Cascade-driven opinion dynamics on social networks

### v1

@@ III-D Polarization
In the context of measuring polarization in social networks, the metrics $NDI$, $GDI$, and $P_{1}$ quantify polarization based on the degree of disagreement among individuals’ opinions. On the other hand, metrics $P_{2}$, $P_{3}$, and $P_{4}$ consider polarization to occur when opinions are significantly distant from a neutral opinion, which is assigned a value of 0.
@@
is referred to as the polarization value and, by definition, is positive if and only if the model is polarizing with respect to the metric $\Phi$. Our previous research has demonstrated that the FJ model is in fact polarizing, and a formula for obtaining polarizing initial opinion vectors for the metrics $P_{2}$, $P_{3}$, and $P_{4}$ have been identified. For what concerns vectors that maximise $P_{2},P_{3}$ polarization, they are correlated to the spectral properties of matrix $H=(I-\Lambda W)^{-1}(I-\Lambda)$ which transforms the initial opinions $\bm{u}$ into the final opinions $\bm{z}$ as described in Equation (2). In particular, we identified $\bm{u}_{B_{2}(1)}$ maximises polarization on the $L_{2}$-ball of radius 1 and $\bm{u}_{B_{2}(t)}=t\cdot\bm{u}_{B_{2}(1)}$ is its scalar multiple that maximises polarization on the $L_{2}$-ball of radius $t$. Furthermore, $\bm{u}_{V_{>1}}$ maximises polarization on the subspace of the specific $H$’s eigenvectors and $\bm{u}_{\max_{P_{2,3}}}$ on the entire domain. For what concerns vectors that maximise $P_{4}$ polarization, we identified $\bm{u}_{B_{1}(1)}$ that maximises polarization on the $L_{1}$-ball of radius 1 and $\bm{u}_{\max_{P_{4}}}$ on the entire domain. Unfortunately, while vectors $\bm{u}_{B_{2}(1)}$, $\bm{u}_{B_{2}(t)}$ and $\bm{u}_{B_{1}(1)}$ can be always found, $\bm{u}_{V_{>1}}$, $\bm{u}_{\max_{P_{2,3}}}$ and $\bm{u}_{\max_{P_{4}}}$ are often impossible to be computed since they are the solutions of NP-hard problems. However, $\bm{u}_{V_{>1}}$ can often be estimated with a heuristic. Please refer to [14] for further details. In Section V we will examine the impact of information cascades when the opinion dynamic process starts with users featuring these polarizing opinions. Notably, we will observe that FJC model induces significantly greater polarization than the FJ for most of the scenarios studied.
@@ IV-A Revisiting the independent cascade model
A post may become obsolete and lose its relevance over time. When this occurs, nodes cease to propagate it, even if some nodes have not been reached. We capture this effect by defining a threshold parameter $L$, known as the interest threshold. Specifically, $L$ represents the maximum hop distance that a post can travel from its source to its destinations. Thus, once a post has reached $L$ hops, its propagation ceases. It is important to note that if this threshold is greater than or equal to the diameter of the network (defined as the longest shortest path between any two nodes in the graph $\mathcal{G}$), all posts will always propagate throughout the entire network, rendering the parameter $L$ ineffective in influencing the resulting cascade.
In our model, an individual identified as the source disseminates a piece of news to the online social network. In the rest of this paper, we streamline our discussion by initially constructing the model under the assumption that each post originates from a single seed node. However, it is worth noting that the model’s flexibility allows for straightforward extension to accommodate multiple sources, akin to the original information cascade model. In fact, including multiple sources simply affects the first level of the information cascade tree, while the rest remains unchanged. We assume (A2) that once the post is shared the first time by the seed node $r$, it spreads outwards like a tree, from the original source to the first, second, third, and subsequent in-neighbours, which we will refer to as the nodes in the first, second, third, … layer. This assumption means that the opinions of parent nodes represent the first exposure a node has in its news feed, thereby solidifying their influence. Following this initial encounter, we maintain that nodes typically become disinterested in posts shared by others. Thus, we will refer to the $l$-th in-neighbourhood as the $l$-th layer or layer $l$ and we will indicate it with $\mathcal{O}_{l}^{(r)}$. Keeping in mind that $L$ is the interest threshold, the entire tree that comprehends all nodes in the aware state of the cascade is denoted as the sequence $\mathcal{O}^{(r)}=\left(\mathcal{O}_{l}^{(r)}\right)_{l=0}^{L}$ of all layers.
A post $p$ generated at node $r$ spreads into the network over the tree $\mathcal{O}^{(r)}$. However, not all nodes in this tree receive the content, because there is a probability $\theta$ that each node will not reshare the post. The subtree of $\mathcal{O}^{(r)}$ that includes only nodes that have received the post is denoted as $\mathcal{T}^{(r,\theta)}=(\mathcal{T}^{(r,\theta)}_{l})_{l=0,\dots,L}$, and it holds that $\mathcal{T}^{(r,\theta)}_{l}\subseteq\mathcal{O}^{(r)}_{l}$ for each $l$. $\mathcal{T}^{(r,\theta)}$ is a stochastic realization of the tree $\mathcal{O}^{(r)}$, as indicated by the superscript $\theta$. In Figure 2, there is a simple example of a network and a corresponding propagation tree.
@@
Figure 2: A toy example of a network in (a) and the corresponding propagation tree for $r=1$ in (b). Nodes $2$ and $3$ update their opinion deterministically, while the others, do it with probability $\theta$.
We can describe it in formal terms in the following way. We refer to the cascade generated by the post $p$ originating in node $r$ as the collection of nodes in the tree $\mathcal{T}^{(r,\theta)}$ and we will denote it with $\mathcal{C}^{(r)}$. Please observe that nodes in layer 1 always belong to the cascade because they receive the post directly from the source, while those of layer $l\geq 2$ belong to the cascade if and only if they receive the post from nodes in layer $l-1$.
@@
$\left\{\begin{array}[]{l}\mathcal{T}^{(r,\theta)}_{0}=\{r\}\\ \mathcal{T}^{(r,\theta)}_{1}=\mathcal{N}^{-}(r)\\ \mathcal{T}^{(r,\theta)}_{l+1}=\left(\bigcup\limits_{\{i\in\mathcal{T}^{(r,\theta)}_{l}:S_{i}=1\}}\mathcal{N}^{-}(i)\right)\setminus\mathcal{T}^{(r,\theta))}_{l}.\end{array}\right.$ (15)
@@
For our convenience, for each node $i\in\mathcal{O}^{(r)}_{l}$ we denote with $\varphi(i)$ the set of nodes who actually influence $i$’s opinion in the stochastic realization of the cascade, i.e the nodes in layer $l-1$ that are in $i$’s out-neighbour and reshare the post. In other terms:
@@ IV-B The cascade-driven Friedkin-Johnsen
We will now discuss how to integrate the independent cascade model described in Section IV-A with the FJ model from Section III-C, resulting in a FJ variant (FJC) where opinions are influenced by the sharing of posts. In this model, opinions are influenced by the sharing of posts among nodes. The FJC model operates iteratively, following specific states and transitions for each node (see Figure 1). At the start of each cascade, every node, except for the designated seed node $r$, begins in the unaware state. During each iteration, nodes $i$ that are exposed to a post from another node $s$ undergo two consecutive actions. First, these exposed nodes transition to the ”aware” state and update their opinions regarding the post. Following this, with a probability $\theta_{s,i}$, these nodes enter the spreader state, sharing the post again if they belong to a layer $l$ that is smaller than the maximum layer $L$. If a node does not meet this condition, the propagation of the post stops, and a new cascade is initiated. We will examine the spreading process of $N$ posts, denoted with $p_{1},\dots,p_{N}$. We make an assumption (A3) that posts propagate one at a time, meaning that simultaneous spread of two posts is not possible. While this assumption is stringent, it aids in designing the model .In future work, we will explore ways to relax this assumption and understand its impact on the model. However, in Section V-C, we demonstrate that the model is robust in real-world settings.
Now, we want to discuss the details of the opinion updating operation. Please observe that, since we want to update opinion layer by layer, we need an asynchronous version of the FJ, similar to that of Equation (3).
We focus on the spreading process of the $n$-th post $p_{n}$ generated at node $r_{n}$ which involves nodes in the cascade $\mathcal{C}^{(r_{n})}$. As the cascade starts, the opinions of nodes are equal to $\bm{u}=\bm{x}(0)$ if $n=1$, i.e. if we are considering the spreading of the first post. Otherwise, they are equal to the final opinion of nodes as the previous post $p_{n-1}$ completed spreading. Every time a node $j$ in the cascade reshares the post $p_{n}$ at a time $t$, it conveys its current opinion $x_{j}(t)$ together with the reposting. A node $i$ in layer $l$ can receive the post (which, as explained above, is a proxy for a node’s opinion) from one or more predecessors $j$. This triggers the update of node $i$’s opinion, which is carried out according to the following update rule:
@@
where $h_{i}$ and $\gamma_{i}$ are defined as in Equations 4 and 5, and $\varphi(i)$ is defined in the Equation (16). The above update rule factors in the average total influence on node $i$ by all its in-neighbours that have reshared the post, hence shared their opinion with $i$. Note that, differently from the asynchrounous FJ update in Equation (3), node $i$ assesses the opinions of its predecessors all at once, not in a pairwise fashion.
@@
1:$\bm{R}$ (vector of post seeds), $\bm{u}$ (initial opinion vector)
2:$\bm{x}$ (final opinion vector)
3:$\bm{x}\leftarrow\bm{u}$triangleright initialize the opinion vector
@@
5:     for $l$ in $\{1,\dots,L\}$ do $\triangleright$ iterate over layers
@@ V Performance Evaluation
In this section, we analyze the performance and validate the FJC model across different social networks. To this end, we conducted two distinct sets of experiments. The first set (Sec. V-B) aimed to compare the opinion dynamics produced by the FJC model with those of the classical FJ model. For this purpose, we used three social networks—two real-world and one synthetic—and generated synthetic cascades on top of them. Specifically, we selected 15 seed nodes per graph to initiate the simulated cascades. The second set of experiments (Sec. V-C) focused on validating the FJC model using real-world cascade data. Here, we evaluated how well the model can approximate opinion evolution based on traces from an online social network, where assumptions (A2) and (A3) do not always hold. Specifically, (A2) posits that cascades spread in a tree-like fashion from the seed node, layer by layer, and (A3) assumes that only one post propagates at a time. Assumption (A1), however, could not be relaxed, as it underpins the model’s formulation of opinion dynamics — it assumes that the content itself is opinion-neutral, and opinions are conveyed solely by the user sharing the post. In the following, we provide a detailed description of the experimental settings and discuss the results obtained.
@@ V-A Common experimental settings
All networks used are unweighted, so the social matrix $\hat{W}$ corresponds to the adjacency matrix. As described in Section III-A, the influence matrix is derived by row-normalizing $\hat{W}$: $w_{ij}=\frac{\hat{w}{ij}}{\sum_{k}\hat{w}{ik}}$. Following our previous work [14], we considered three strategies for setting node susceptibility. The first two rely on PageRank centrality, a measure that effectively captures node influence [23]. For each node $i$ with PageRank score $\mathrm{P}_{i}$, we assigned $\lambda_{i}$ either the rescaled value of $\mathrm{P}_{i}$ or $\mathrm{P}_{i}^{-1}$, both mapped to the interval $(0,1)$, thus giving higher susceptibility to more (or less) central nodes. In the third strategy, all nodes were assigned a fixed susceptibility value of 0.6. These variations allow us to explore how susceptibility interacts with node centrality. Finally, since our primary goal is to study the impact of cascades on polarization, we initialized opinions using the polarizing vectors defined in Section III-D, ensuring a consistent and controlled setup for measuring polarization dynamics.
V-B FJC vs FJ – with synthetic cascades

### v2

@@ III-D Polarization
In the context of measuring polarization in social networks, the metrics $NDI$, GDIGDI, and $P_{1}$ quantify polarization based on the degree of disagreement among individuals’ opinions. On the other hand, metrics $P_{2}$, $P_{3}$, and $P_{4}$ consider polarization to occur when opinions are significantly distant from a neutral opinion, which is assigned a value of 0.
@@
is referred to as the polarization value and, by definition, is positive if and only if the model is polarizing with respect to the metric $\Phi$. Our previous research [7] showed that the FJ model is inherently polarizing and provided formulas to derive polarizing vectors for the polarization metrics $P_{2}$, $P_{3}$, and $P_{4}$. These vectors are closely linked to the spectral properties of the matrix $H=(I-\Lambda W)^{-1}(I-\Lambda)$, which maps initial opinions $\bm{u}$ to final opinions $\bm{z}$ as in Equation (2). Depending on the considered sub-domain of the opinion space, different polarizing vectors can be identified. In this paper, for $P_{4}$ polarization we use $\bm{u}_{B_{1}(1)}$, while for $P_{2}$ and $P_{3}$ polarization we rely on $\bm{u}_{B_{2}(1)}$, $\bm{u}_{B_{2}(t)}$, and the heuristic $\bm{u}_{V_{>1}}^{heu}$, listed in order of increasing polarization strength. For a detailed mathematical treatment and derivations, we refer the reader to [7]. In Section V we will examine the impact of information cascades when the opinion dynamic process starts with users featuring these polarizing opinions. Notably, we will observe that FJC model induces significantly greater polarization than the FJ for most of the scenarios studied.
@@ IV-A Revisiting the independent cascade model
In our model, an individual identified as the source disseminates a piece of news to the online social network. In the rest of this paper, we streamline our discussion by initially constructing the model under the assumption that each post originates from a single seed node. However, it is worth noting that the model’s flexibility allows for straightforward extension to accommodate multiple sources, akin to the original information cascade model. In fact, including multiple sources simply affects the first level of the information cascade tree, while the rest remains unchanged. We assume (A2) that once the post is shared the first time by the seed node $r$, it spreads outwards like a tree, from the original source to the first, second, third, and subsequent in-neighbours, which we will refer to as the nodes in the first, second, third, … layer. This assumption means that the opinions of parent nodes represent the first exposure a node has in its news feed, thereby solidifying their influence. Following this initial encounter, we maintain that nodes typically become disinterested in posts shared by others. Thus, we refer to the $l$-th in-neighbourhood as the $l$-th layer (or layer $l$), and denote it by $\mathcal{O}_{l}^{(r)}$. The depth of the graph is given by the eccentricity of the root node, denoted by $\epsilon^{(r)}$, which corresponds to the maximum distance from $r$ to any other node. Accordingly, the entire tree comprising all nodes in the aware state of the cascade is represented as the sequence $\mathcal{O}^{(r)}=\left(\mathcal{O}l^{(r)}\right)_{l=0}^{\epsilon^{(r)}}$ of all layers.
A post $p$ generated at node $r$ spreads into the network over the tree $\mathcal{O}^{(r)}$. However, not all nodes in this tree receive the content, because there is a probability $\theta$ that each node will not reshare the post. The subtree of $\mathcal{O}^{(r)}$ that includes only nodes that have received the post is denoted as $\mathcal{T}^{(r,\theta)}=(\mathcal{T}^{(r,\theta)}_{l})_{l=0,\dots,\epsilon^{(r)}}$, and it holds that $\mathcal{T}^{(r,\theta)}_{l}\subseteq\mathcal{O}^{(r)}_{l}$ for each $l$. $\mathcal{T}^{(r,\theta)}$ is a stochastic realization of the tree $\mathcal{O}^{(r)}$, as indicated by the superscript $\theta$. In Figure 2, there is a simple example of a network and a corresponding propagation tree.
@@
Figure 2: A toy example of a network in (a) and the corresponding propagation tree for $r=1$ in (b). Nodes 22 and $3$ update their opinion deterministically, while the others, do it with probability $\theta$.
We can describe it in formal terms in the following way. We refer to the cascade generated by the post $p$ originating in node $r$ as the collection of nodes in the tree $\mathcal{T}^{(r,\theta)}$ and we will denote it with $\mathcal{C}^{(r)}$. Please observe that nodes in layer 1 always belong to the cascade because they receive the post directly from the source, while those of layer $l\geq 2$ belong to the cascade if and only if they receive the post from nodes in layer l1l-1.
@@
[…]
@@
For our convenience, for each node $i\in\mathcal{O}^{(r)}_{l}$ we denote with $\varphi(i)$ the set of nodes who actually influence $i$’s opinion in the stochastic realization of the cascade, i.e the nodes in layer l1l-1 that are in $i$’s out-neighbour and reshare the post. In other terms:
@@ IV-B The cascade-driven Friedkin-Johnsen
We will now discuss how to integrate the independent cascade model described in Section IV-A with the FJ model from Section III-C, resulting in a FJ variant (FJC) where opinions are influenced by the sharing of posts. The FJC model operates iteratively, following specific states and transitions for each node (see Figure 1). At the start of each cascade, every node, except for the designated seed node $r$, begins in the unaware state. During each iteration, nodes $i$ that are exposed to a post from another node $s$ undergo two consecutive actions. First, these exposed nodes transition to the ”aware” state and update their opinions regarding the post. Following this, with a probability $\theta_{s,i}$, these nodes enter the spreader state, sharing the post again if they belong to a layer $l$ that is smaller than the maximum layer $\epsilon^{(r)}$. If a node does not meet this condition, the propagation of the post stops, and a new cascade is initiated. We will examine the spreading process of $N$ posts, denoted with $p_{1},\dots,p_{N}$. We make an assumption (A3) that posts propagate one at a time, meaning that simultaneous spread of two posts is not possible. While this assumption is stringent, it aids in designing the model .In future work, we will explore ways to relax this assumption and understand its impact on the model. In Section V-B, we discuss the model effectiveness in real-world settings.
Now, we want to discuss the details of the opinion updating operation. Please observe that, since we want to update opinion layer by layer, we need an asynchronous version of the FJ, similar to that of Equation (3). We focus on the spreading process of the $n$-th post $p_{n}$ generated at node $r_{n}$ which involves nodes in the cascade $\mathcal{C}^{(r_{n})}$. As the cascade starts, the opinions of nodes are equal to $\bm{u}=\bm{x}(0)$ if $n=1$, i.e. if we are considering the spreading of the first post. Otherwise, they are equal to the final opinion of nodes as the previous post $p_{n-1}$ completed spreading. Every time a node $j$ in the cascade reshares the post $p_{n}$ at a time $t$, it conveys its current opinion $x_{j}(t)$ together with the reposting. A node $i$ in layer $l$ can receive the post (which, as explained above, is a proxy for a node’s opinion) from one or more predecessors $j$. This triggers the update of node $i$’s opinion, which is carried out according to the following update rule:
@@
where $h_{i}$ and $\gamma_{i}$ are defined as in Equations 4 and 5, and $\varphi(i)$ is defined in the Equation (16). The above update rule factors in the average total influence on node $i$ by all its in-neighbours that have reshared the post in the previous layer. This means that the average is calculated only from those neighbors who have already engaged with the post and shared their opinions, thereby influencing the opinion of node $i$. Note that, differently from the asynchrounous FJ update in Equation (3), node $i$ assesses the opinions of its predecessors all at once, not in a pairwise fashion.
@@
[…]
@@
5:  for $l$ in $\{1,\dots,\epsilon^{(r)}\}$ do $\triangleright$ iterate over layers
@@ V Performance Evaluation
In this section, we analyze the performance and validate the FJC model across different social networks. To this end, we conducted two distinct sets of experiments. The first set of experiments (Sec. V-B) focused on validating the FJC model using real-world cascade data. Here, we evaluated how well the model can approximate opinion evolution based on traces from an online social network, where assumptions (A2) and (A3) do not always hold. Specifically, (A2) posits that cascades spread in a tree-like fashion from the seed node, layer by layer, and (A3) assumes that only one post propagates at a time. Assumption (A1), however, could not be relaxed, as it underpins the model’s formulation of opinion dynamics — it assumes that the content itself is opinion-neutral, and opinions are conveyed solely by the user sharing the post. In the following, we provide a detailed description of the experimental settings and discuss the results obtained. The second set (Sec. V-C) aimed to compare the opinion dynamics produced by the FJC model with those of the classical FJ model. For this purpose, we used three social networks—two real-world and one synthetic—and generated synthetic cascades on top of them. Specifically, we selected 15 seed nodes per graph to initiate the simulated cascades.
@@ V-A Common experimental settings
All networks used are unweighted, so the social matrix $\hat{W}$ corresponds to the adjacency matrix. As described in Section III-A, the influence matrix is derived by row-normalizing $\hat{W}$: $w_{ij}=\frac{\hat{w}{ij}}{\sum_{k}\hat{w}{ik}}$. Following our previous work [7], we considered three strategies for setting node susceptibility. The first two rely on PageRank centrality, a measure that effectively captures node influence [25]. For each node $i$ with PageRank score $\mathrm{P}_{i}$, we assigned $\lambda_{i}$ either the rescaled value of $\mathrm{P}_{i}$ or $\mathrm{P}_{i}^{-1}$, both mapped to the interval $(0,1)$, thus giving higher susceptibility to more (or less) central nodes. In the third strategy, all nodes were assigned a fixed susceptibility value of 0.6. These variations allow us to explore how susceptibility interacts with node centrality. Finally, since our primary goal is to study the impact of cascades on polarization, we initialized opinions using the polarizing vectors defined in Section III-D, ensuring a consistent and controlled setup for measuring polarization dynamics. Final opinions under cascades were then compared with those of the FJ model, whose outcomes are analytically derived via Equation (2) using the same parameters, except for the reposting probability $\theta$, which is not applicable to FJ.
V-B FJC with real cascades
@@ V-B1 Dataset description and experimental setting
We validate the model using a real-world dataset from Bhowmick et al.  [46], which contains Twitter (now X) data related to the 2015 Nepal earthquake. This dataset is well-suited for our analysis as it includes both the follower network and tweet/retweet histories. Unlike the synthetic networks in Sec. V-C, this is a directed graph, with multiple connected components. Since FJ-based models operate within a connected structure, we extract the largest strongly connected component (LSCC). Some cascades span disconnected components, suggesting hidden links, so we restrict our analysis to cascades occurring entirely within the LSCC. The resulting graph contains 220,974 nodes, 16,772,693 edges, and 32,313 cascades. Initial opinions are set using the polarizing vectors described in Section III-D. However, their computation requires to compute of the eigensystem of an $n\times n$ dense matrix, which becomes impractical due to the extensive size of the network, even when employing optimized algorithms. To tackle this challenge, we derived a subgraph known as Nepal top cascades, which includes nodes from the 500 longest cascading events. The rationale behind focusing on the longest cascades is that they inherently present a greater challenge for our model, as they involve more complex and prolonged interactions among nodes and increase the likelihood that cascades can overlap (thus, violating assumption (A3)). Specifically, we ensured that all nodes participating in these cascades are included in the subgraph using a heuristic for the directed Steiner Tree problem [47]. Table I summarizes the key properties of this subgraph, which encompasses over 6,000 cascades.
TABLE I: Summary of the real-world Nepal dataset subgraphs used for validation
#nodes #edges #cascades max cascade length
Nepal top cascades 26321 4222064 6083 21
The dataset $D$ provides a sequence of tuples $\mathcal{C}_{D}=\big{(}(t_{1},p_{1},i_{1}),(t_{2},p_{2},i_{2}),\dots\big{)}$, where node $i_{k}$ shares post $p_{k}$ at time $t_{k}$, representing the temporal evolution of cascades. We assume that this sequence of reshares drives the opinion dynamics in the real world: each reshare is treated as an expression of opinion. At time $t_{k}$, when node $i_{k}$ shares a post, its followers update their opinions according to Equation (17). This process relies solely on assumption (A1); assumptions (A2) and (A3) are not used. We refer to this approach as real opinion dynamics. It is a data-driven heuristic and cannot be analyzed analytically.
Our goal is to evaluate how well this real dynamic is captured by the proposed FJC model (we also include the classical FJ model for reference). For the FJC model, we need to estimate the resharing probability $\theta_{i}$ for each node. Inspired by weighted cascade models [36, 48], we define $\theta_{i}$ as the ratio between the number of posts reshared by node $i$ and the number of posts it has seen from its followees. Table II shows the distribution of $\theta_{i}$, which is generally low, indicating that most nodes rarely reshare posts. Please note that these values are consistent with the retweet probabilities in [49], remaining below 0.02. Unlike the real dynamics, the FJC model processes cascades in a tree manner and sequentially. Specifically, posts are propagated from the sources to the network layer-by-layer, in accordance with assumption (A1) and each cascade starts only after the previous one ends, in accordance with assumption (A3).
TABLE II: Statistics of followers, followed and $\theta$ in the Nepal subgraph
Nepal top cascades
Min. 1st Qu. Median 3rd Qu. Max. Mean N.A.
Followers 1.0 23.0 59.0 148.0 11022.0 160.4 -
Followees 1.0 53.0 109.0 204.0 6851.0 160.4 -
$\theta$ 0.0000 0.0000 0.0000 0.0000 1.0000 0.0156 681
V-B2 Opinion evolution of real opinion dynamics vs FJC and FJ
Due to the large size of the networks, we cannot visualize the polarizing vectors, since the plots would be unreadable. Instead, our main objective with these datasets is to evaluate the robustness of the FJC model against the real opinion dynamics. We do so by comparing the distributions of opinion shifts—i.e., the differences between final and initial opinions—for the real evolution (labeled Real) and the predictions of FJC and FJ. As shown in Figure 3, FJC closely mirrors the real dynamics observed in the Nepal top cascades dataset, outperforming the standard FJ model. Please note that long cascades are more likely to violate assumption (A3), reducing FJC’s accuracy. Nevertheless, FJC remains surprisingly reliable even in this more challenging scenario.
Figure 3: ECDF of the opinion shift of nodes, i.e. final opinion minus initial opinion, for the Nepal top cascades in the three different opinion dynamics. (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
V-B3 Polarization
Figure 4 illustrates the polarization outcomes across different settings. As expected, the Real opinion dynamics remain almost entirely null, since opinion shifts are negligible in all scenarios (Figure 3). Similarly, the cascade-based dynamics (FJC) exhibit very low—or even slightly negative—polarization. Both stand in sharp contrast to the classical FJ model, which consistently produces positive polarization with substantially higher magnitude. This behavior can be explained by Table II, which shows that most nodes have $\theta$ values close to zero, indicating very limited reposting activity. Consequently, the opinion dynamics remain largely inactive, preventing polarization from emerging. In Section V-C, we will further validate this hypothesis on synthetic networks, also exploring scenarios with larger $\theta$ values to assess their impact on polarization.
V-C FJC vs FJ – with synthetic cascades
V-C1 Dataset description and experimental setting